What is the sum of the solutions of 2|x-1|-4=-2
a. 1 1/2
b. 2
c. 2
d. 0

Found 2 solutions by ewatrrr, Edwin McCravy:
Answer by ewatrrr(24785)

 
Hi,
2|x-1|-4=-2
0r
|x-1| = 1
x - 1 = 1, x = 2
x - 1 = -1, x = 0
2 + 0 = 2

Answer by Edwin McCravy(20055)

To avoid doing people's homework for them, I'll do one
that is solved exactly the same way.  Use it as a model
for solving yours:

3|x-2| - 8 = -2

Add 8 to both sides

    3|x-2| = 6

Divide both sides by 3

     |x-2| = 2

Make two equations:

x-2 = 2    OR    x-2 = -2
  x = 4    OR      x = 0

The sum of the two solutions is 4+0 or 4.

Edwin